What is the mass of exactly 1 mole of each of the compounds? NO, Fe(CIO): Fe(CIO),
Answer #1

Solution : As we know that mass of the 1mole of the compound is called molar mass. Ans molar mass of a compound can be calculated as

N2O4=2 * mass of N + 4 * mass of O

= 2 * 14 +4 * 16 = 28 +64 = 92 g

Fe(ClO)2= mass of Fe +2 * (mass of Cl +mass of O)

=56+2 * (35.5+ 16)

= 56+ 2 * 51.5

= 56 + 103.0=159 g

Fe(ClO4)2 =mass of Fe + 2 * (mass of Cl + 4 * mass of O)

=56 +2 * ( 35.5 + 4 * 16)

= 56 +2 * (35.5 + 64)

=56 + 2 * 99.5 = 56 + 199.0 = 255.0 g